# Advances in Radiance Field for Dynamic Scene: From Neural Field to Gaussian Field

### v1

@@ 4.3 Reconstructing with Non-rigid Motion
4D Spacetime. A complete light field is traditionally represented as a 7D plenoptic function $f(\mathbf{x},\mathbf{d},\lambda,t)$ [190], which simplifies to a 5D function for static scenes by removing time and wavelength dependencies. For dynamic scenes, researchers reintroduce time $t$, creating a 6D plenoptic function that models the scene as a continuous 4D spatiotemporal volume [26, 91]. This unified representation implicitly encodes scene geometry, appearance, and motion within the same field structure, with properties like density and color varying temporally without explicitly modeling motion vectors. To enhance efficiency and visual fidelity, recent methods decompose scenes into time-invariant static backgrounds (5D) and time-dependent [180, 94] or latent code conditioned [41] dynamic foregrounds (6D), blended via learned weights and often guided by segmentation masks [79, 191] or self-supervised techniques [192]. Implementation approaches have evolved from MLPs to more efficient structures: some methods use 4D neural voxels for accelerated rendering [90, 193, 194], while others leverage 3D Gaussian splatting to model 4D spacetime as sequential slices of 3D space with time-dependent Gaussian properties [27, 181], enabling real-time rendering of complex non-rigid motion while maintaining temporal consistency.
Canonical Space with Deformation Field. In contrast to frame-by-frame 4D spacetime optimization that entangles motion and scene representation within a single radiance field, an alternative approach decomposes dynamic scenes into a static reference volume (canonical space) and its temporal evolution pattern (deformation field). The canonical space serves as a consistent template representation, while the deformation field describes how each point moves across time. For structured subjects like human bodies or faces, the canonical space often represents a neutral pose or expression. For general scenes, it typically corresponds to the scene configuration at an initial time point $t_{0}$.
@@
In neural radiance field implementations, this approach typically employs backward deformation—points sampled along camera rays in the observation space are transformed back to canonical space through the deformation field, as their density and color properties are stored in the canonical representation. In contrast, explicit representations like 3D Gaussian fields can directly apply forward deformation, warping each Gaussian primitive from canonical to observation space. Neither approach alone guarantees perfect consistency between spaces, leading some methods to implement bijective deformation fields that maintain correspondences in both directions.
@@
For extended sequences with substantial motion or appearance changes, a single global canonical space often proves insufficient. In such cases, multiple local canonical spaces (keyframes) shared by temporal subwindows provide a more effective solution [184]. This approach allows nearby frames to reference the same canonical space while temporally distant frames leverage different reference volumes, better accommodating dramatic transformations while maintaining local consistency.
@@
Frame-to-Frame Flow Field. An alternative approach to representing dynamic scenes involves modeling motion as point correspondences between consecutive frames, known as frame-to-frame motion fields or scene flow fields. These fields typically formulate the motion relationship between adjacent frames, where smaller displacements make the motion patterns easier to learn and model. This flow field is generally implemented as a 4D function that maps 3D spatial positions and a 1D time parameter to corresponding displacement vectors [38, 37, 195].
@@
Rather than representing flow in only one direction, bidirectional approaches enhance reconstruction quality. For example, Li et al. [13] utilize both forward and backward flow fields within the same framework, establishing point correspondences between frames $i$ and $j$ along rays. When points from frame $i$ move to frame $j$ along flow field $f_{i\rightarrow j}$, the rendered results should maintain consistency with frame $j$, and vice versa. This bidirectional consistency effectively enables information sharing between adjacent frames, serving as a powerful constraint that enhances learning efficiency [96].
@@
Beyond direct flow field learning, an alternative formulation treats flow fields as the integration of velocity fields over time [96, 38, 19, 19]. While flow fields are typically constrained by smoothness and continuity regularizations, velocity fields offer additional physical constraints and directional information through their vector nature. Li et al. [196] demonstrate this by incorporating physical laws as supervision through physics-informed neural networks (PINNs), enabling applications like future frame extrapolation, motion transfer, and semantic decomposition. Regardless of whether flow fields or velocity fields are employed, 2D optical flow provides valuable supervision signals for learning [38, 185, 13, 186], particularly in monocular settings where depth information is limited.
@@
Point Tracking: Traditional approaches to dynamic scene reconstruction employ either canonical spaces with deformation fields or frame-to-frame progressive evolution. Both methods calculate point correspondences indirectly through proxy spaces (canonical or reference frames), which becomes problematic for long-term correspondence as errors accumulate along transformation chains.
@@
An effective alternative models each point’s movement as a continuous trajectory, a time-dependent function that directly describes position at any moment in the continuous spacetime domain [181, 23]. This global trajectory formulation represents the complete motion path as a unified function, eliminating cascading errors from sequential transformations. With this approach, a point’s geometry and appearance can be modeled as time-varying radiance properties while maintaining temporal consistency [18].
@@
For multi-view capture scenarios, 3D Gaussian Splatting methods can initialize the scene representation from the first frame and subsequently track each Gaussian primitive’s movement through space over time [45, 17]. This tracking approach can maintain consistent appearance properties (density and color) while updating positions, providing a more robust foundation for dynamic scene reconstruction with greater temporal coherence.
Factorization: The concept of factorization for radiance fields emerged from static scene reconstruction, where methods like TensoRF [5] decompose 3D space into low-rank 1D vectors or 2D matrices to improve computational and memory efficiency. This approach has been successfully extended to 4D dynamic scenes through spatiotemporal factorization.
@@
Hexplane [15] decomposes the 4D domain into six feature planes, where point features are sampled via interpolation and concatenated to predict density and color. Similarly, K-planes [14] offers a unified approach for both static and dynamic scenes—factorizing static 3D space into $xy,yz,xz$ planes while representing dynamic 4D spacetime with $xt,yt,zt$ planes, incorporating multi-scale sampling for enhanced representation. This efficient spatiotemporal factorization enables higher grid resolution and rapid convergence, making it a foundation for numerous subsequent methods [198, 42, 199, 200]. For instance, 4K4D [152] extends K-plane’s 4D feature grid factorization to achieve real-time performance at 4K resolution, while Wu et al. [11] use factorized spatiotemporal structures to encode per-Gaussian features for deformation field decoding.
Beyond basic spatiotemporal factorization, motion fields can also be factorized into a compact set of basis functions. This approach represents complex motion using a few representative spatial deformation patterns, enhancing temporal coherence while improving learning stability and storage efficiency. Li et al. [20] model the motion field as spatially decomposed motion bases with time-varying coefficients, using pixel-aligned features sampled from source to target views and fused by a ray transformer. While some methods represent both motion bases and coefficients as learnable neural network parameters [201, 202], others leverage sinusoidal bases [203, 23, 188].
@@
For Gaussian-based representations, Kratimenos et al. [189] factorize 3D Gaussian motion into a small number of motion bases—significantly fewer than the number of Gaussian primitives—with regularization applied to the motion coefficients to ensure plausible movements. Das et al. [187] propose a two-stage approach: first learning a coarse proxy using factorized motion bases and low-rank coefficients, then initializing local volumes with 3D Gaussians refined through adaptive densification. The shared motion bases in the coarse stage force information sharing between timesteps, providing essential regularization for sparsely observed dynamic regions.
@@ 4.4 Reconstructing with Hybrid Motion
Dynamic objects in real-life scenarios often exhibit complex motion patterns that combine rigid, articulated, and non-rigid components. In autonomous driving scenes, for example, vehicles primarily undergo rigid motion, while pedestrians and cyclists move in more complex, non-rigid ways. To address this diversity, Fischer et.al [51] leveraged tracking results with bounding boxes to represent rigid vehicles while employing non-rigid motion fields to model other dynamic objects. Taking a more comprehensive approach, OmniRe [80] developed a framework that combines distinct motion representations within a dynamic scene graph: rigid nodes for vehicles, SMPL nodes for articulated pedestrian motion, and deformable nodes for general non-rigid objects—creating a unified framework for urban scene reconstruction.
@@ 4.5.1 Auxiliary Information
Depth Information serves as a crucial geometric cue for 3D scene reconstruction, particularly valuable in dynamic scenarios where it helps mitigate ambiguities in scale, motion, and geometry. This spatial information can be acquired through various approaches, ranging from direct sensor measurements via RGB-D cameras or LiDAR systems to algorithmic estimation techniques. In neural radiance fields, depth is rendered by replacing the color value in the volume rendering equation (Eq. 3), while in 3D Gaussian splatting, depth is obtained by rasterizing the Gaussian positions during splatting (Eq. 4). Through gradient backpropagation, supervising the rendered depth in observation space significantly aids the learning of both radiance fields and motion fields [94, 204], particularly beneficial in challenging outdoor environments [108, 106, 111, 107]. When dedicated depth sensors are unavailable, monocular depth estimation methods [205, 206] can provide valuable geometric cues despite scale ambiguity [16, 182, 93, 204]. Beyond providing additional geometric supervision, depth information enables importance sampling in regions near object surface, substantially reducing unnecessary computational overhead in volume-based rendering methods [207, 79].
@@
Surface Normals capture fine-grained geometric details that might be missed in continuous and smooth neural representations, making them valuable for high-fidelity reconstruction. In dynamic scenes, surface normals are particularly useful in tracking shape deformations [85, 157] and modeling view-dependent reflections [208]. Recent research has successfully leveraged normal information to recover detailed human surface geometry [209, 210, 211], and enhance scene reconstruction [212, 213]. Surface normals are computationally convenient to obtain through automatic differentiation frameworks like PyTorch and TensorFlow, where they can be calculated by differentiating depth with respect to spatial coordinates in volume rendering approaches [117, 214]. In 3D Gaussian splatting, normals can be rendered by taking the direction of the shortest axis of the anisotropic Gaussians [157]. For supervision purposes, pseudo ground truth normals can be derived by being directly estimated via foundation models like Metric3D [215] or calculated from template meshes [36, 159, 167].
@@
Semantic Information serves as a powerful auxiliary cue for dynamic scene reconstruction, enabling more nuanced understanding and accurate modeling of moving objects. In dynamic scenes, object-level semantic segmentation provides valuable silhouettes or foreground masks that help localize dynamic objects and decompose scenes into static backgrounds and dynamic foregrounds [216, 208, 94]. This semantic decomposition is particularly valuable for complex urban environments and scenes with multiple moving entities [22, 106]. Beyond object-level segmentation, part-level semantic information recognizes the rigid components of articulated objects [217], enforcing part consistency throughout the dynamic reconstruction process and enabling more detailed part-wise reconstruction [218]. At the finest granularity, pixel-level semantic features facilitate robust point tracking and establish reliable correspondences across frames [219, 131, 186, 220, 132], which is essential for tracking deformations and understanding complex motion patterns.
@@
Data-driven Priors significantly enhance dynamic scene reconstruction by leveraging implicit knowledge from large-scale datasets to provide valuable constraints during optimization. Foundation models trained on diverse data deliver robust priors across multiple modalities, including Depth Anything [206] for monocular depth estimation, Metric3D [215] for surface normals, and Segment Anything [221] for semantic segmentation—all serving as effective supervision signals without specialized sensors. For motion understanding, optical flow models like UniMatch [222] establish dense correspondences between frames [86, 161, 162, 223], while object tracking models generate reliable trajectories for rigid objects and statistical parametric models such as SMPL [33] and SMPL-X [82] encode strong priors about human body shape and pose. Additionally, visual foundation models like DINO [104] extract semantic features that maintain consistency across different viewpoints and lighting conditions, providing robust cues for establishing correspondences in challenging scenarios [195, 165, 130, 177], ultimately improving reconstruction quality in situations with sparse views, fast motion, or challenging lighting.
@@ 4.5.2 Regularization
In real-world scenarios, object motion exhibits continuity–neighboring points tend to move similarly, and trajectories evolve gradually over time. To enforce this smoothness mathematically, Total Variation (TV) loss encourages motion fields $f$ to be piecewise constant, creating natural transitions between differently moving regions [91, 14, 224]:
$\mathcal{L}_{TV}=\sum_{i,j}\left|f(i+1,j)-f(i,j)\right|+\left|f(i,j+1)-f(i,j)\right|.$ (24)
Complementing TV loss, Laplacian loss penalizes sharp gradient changes, ensuring smoother transitions in deformation fields by operating on second-order derivatives [153].
Beyond smoothness, maintaining the physical structure of objects during deformation is crucial for realism. The As-Rigid-As-Possible (ARAP) constraint preserves local structural integrity by maintaining relative distances between neighboring points. This approach is particularly effective for short-term local rigidity, ensuring that each small neighborhood deforms as rigidly as possible. While ARAP focuses on local rigidity, isometric loss ensures global structural consistency by preserving geodesic distances across the entire shape:
$\mathcal{L}_{iso}=\sum_{i=1}^{N_{p}}\sum_{j\in\mathcal{N}_{k}(i)}\left\|d\left(\mathbf{x}_{i}^{\mathcal{t_{0}}},\mathbf{x}_{j}^{\mathcal{t_{0}}}\right)-d\left(\mathbf{x}_{i}^{t},\mathbf{x}_{j}^{t}\right)\right\|_{1},$ (25)
In addition to these, the divergence loss encourages volume preservation by constraining deformation fields to primarily consist of translations and rotations. This property is critical for realistic modeling of real-world object deformations where material volume typically remains constant during movement.
Besides these physical constraints, scene priors contribute significantly to guiding reconstruction. For example, applying $L_{1}$ regularization on motion fields biases the scene toward static components–a reasonable assumption given that dynamic objects typically occupy only a small portion of the scene [26, 192]. This regularization not only favors a mostly static background but also facilitates a more robust static-dynamic decomposition during analysis. Moreover, cycle consistency constraints play a critical role in verifying reconstruction accuracy. By enforcing consistent correspondences between features, motions, or colors across frames, these constraints help maintain temporal coherence [38, 13, 195]. However, proper handling of occlusions is crucial here, as consistency can only be enforced for points that remain visible in multiple frames. Finally, for volumetric radiance field representations, opacity regularization is employed to promote binary outcomes (0 or 1). This approach mitigates artifacts such as floating regions by reinforcing the assumption that, in most cases, objects are either fully opaque or absent [26, 192].
Together, these complementary regularization strategies guide the optimization process, ensuring that the resulting dynamic scene reconstructions are physically accurate, temporally consistent, and visually plausible.

### v2

@@ 4.3 Reconstructing with Non-rigid Motion
4D Spacetime is a unified representation that implicitly encodes scene geometry, appearance, and motion within a single radiance field, with properties like density and color varying temporally without explicitly modeling motion vectors [26, 81]. To enhance efficiency and visual fidelity, recent methods decompose scenes into time-invariant static backgrounds (5D) and time-dependent [166, 84] or latent code conditioned [41] dynamic foregrounds (6D), blended via learned weights and often guided by segmentation masks [71, 182] or self-supervised techniques [183]. Implementation approaches have evolved from MLPs to more efficient structures: some methods use 4D neural voxels for accelerated rendering [80, 184, 185], while others leverage 3D Gaussian splatting to model 4D spacetime as sequential slices of 3D space with time-dependent Gaussian properties [27, 167], enabling real-time rendering of complex non-rigid motions.
@@
Canonical Space with Deformation Field approach decomposes dynamic scenes into a static reference volume (canonical space) and its temporal evolution pattern (deformation field) [186, 9, 169, 187]. In neural radiance field implementations, this approach typically employs backward deformation to transform sampled points along camera rays in the observation space back to canonical space [9, 8, 10], as their density and color properties are stored in the canonical frame. In contrast, explicit representations like 3D Gaussian fields can directly apply forward deformation, warping each Gaussian primitive from canonical to observation space [12, 83]. Neither approach alone guarantees perfect consistency between spaces, leading some methods to implement bijective deformation fields that maintain correspondences in both directions [84].
@@
For extended sequences with substantial motion or appearance changes, a single global canonical space often proves insufficient. In such cases, multiple local canonical spaces (keyframes) shared by temporal subwindows provide a more effective solution [170]. This approach allows nearby frames to reference the same keyframe while temporally distant frames leverage different keyframes, better accommodating dramatic transformations while maintaining local consistency.
@@
Frame-to-Frame Flow Field models dynamic motion as point correspondences between consecutive frames, known as frame-to-frame scene flow fields. These fields typically formulate the motion relationship between adjacent frames, where smaller displacements make the motion patterns easier to learn and model. This flow field is generally implemented as a 4D function that maps 3D spatial positions and a 1D time parameter to corresponding displacement vectors [38, 37, 188].
@@
Rather than representing flow in only one direction, bidirectional approaches enhance reconstruction quality. For example, Li et al. [13] utilize both forward and backward flow fields within the same framework, establishing point correspondences between frames $i$ and $j$. When points from frame $i$ move to frame $j$ along flow field $f_{i\rightarrow j}$, the rendered results should maintain consistency with frame $j$, and vice versa. This bidirectional consistency effectively enables information sharing between adjacent frames, serving as a powerful constraint that enhances learning efficiency [85].
@@
Beyond direct flow field learning, an alternative formulation treats flow fields as the integration of velocity fields over time [85, 38, 19]. While flow fields are typically constrained by smoothness and continuity regularizations, velocity fields offer additional physical constraints and directional information through their vector nature. Li et al. [189] demonstrate this by incorporating physical laws as supervision through physics-informed neural networks (PINNs), enabling applications like future frame extrapolation, motion transfer, and semantic decomposition. Regardless of whether flow fields or velocity fields are employed, 2D optical flow provides valuable supervision signals for learning [38, 171, 13, 172], particularly in monocular settings where depth information is limited.
@@
Point Tracking models each point’s movement as a continuous trajectory, a time-dependent function that directly describes position at any moment in the continuous spacetime domain [167, 23]. This global trajectory formulation represents the complete motion path as a time-modulated function, eliminating accumulated errors from sequential transformations, like Eq. 13 in canonical spaces or Eq. 17 in frame-to-frame flow field. With this approach, a point’s geometry and appearance can be modeled as time-varying parameters while maintaining temporal consistency [18].
@@
For multi-view capture scenarios, methods based on 3DGS can initialize the scene representation from the first frame and subsequently track each Gaussian primitive’s movement through space over time [45, 17]. This tracking approach can maintain consistent properties like opacity and color while updating positions, rotation, and scale, providing a more robust foundation for dynamic scene reconstruction with greater temporal coherence.
@@
Factorization emerged from static scene reconstruction and had been successfully extended to 4D dynamic scenes through hyperplane-based factorization. Hexplane [15] decomposes the 4D domain into six feature planes, where point features are sampled via interpolation and concatenated to predict density and color. Similarly, K-planes [14] offers a unified approach for both static and dynamic scenes—factorizing static 3D space into $xy,yz,xz$ planes while representing dynamic 4D spacetime with $xt,yt,zt$ planes, incorporating multi-scale sampling for enhanced representation. This efficient factorization enables higher grid resolution and rapid convergence, making it a foundation for numerous subsequent methods [191, 42, 192, 193]. For instance, 4K4D [142] extends K-plane’s 4D feature grid factorization to achieve real-time performance at 4K resolution, while Wu et al. [11] use factorized grid planes to encode per-Gaussian features for deformation field decoding.
@@
Beyond hyperplane-based factorization, basis-driven decomposition represents complex motion using a few representative spatial deformation patterns, enhancing temporal coherence while improving learning stability and storage efficiency. Li et al. [20] model the motion field as spatially decomposed motion bases with time-varying coefficients, using pixel-aligned features sampled from source to target views and fused by a ray transformer. While some methods represent both motion bases and coefficients as learnable neural network parameters [194, 195], others leverage sinusoidal bases [196, 23, 174]. For Gaussian-based representations, Kratimenos et al. [175] factorize 3D Gaussian motion into a small number of motion bases–significantly fewer than the number of Gaussian primitives–with regularization applied to the motion coefficients to ensure plausible movements. Das et al. [173] propose a two-stage approach: first learning a coarse proxy using factorized motion bases and low-rank coefficients, then initializing local volumes with 3D Gaussians refined through adaptive densification. The shared motion bases in the coarse stage force information sharing between timesteps, providing essential regularization for sparsely observed dynamic regions.
@@ 4.4 Reconstructing with Hybrid Motion
In autonomous driving scenes, vehicles primarily undergo rigid motion, while pedestrians and cyclists move in more complex, non-rigid ways. To address this diversity, Fischer et. al [51] leveraged tracking results with bounding boxes to represent rigid vehicles while employing non-rigid motion fields to model other dynamic objects. Taking a more comprehensive approach, OmniRe [72] developed a framework that combines distinct motion representations within a dynamic scene graph: rigid nodes for vehicles, SMPL nodes for articulated pedestrian, and deformable nodes for general non-rigid objects, creating a more complete representation for urban scene reconstruction.
@@ 4.5.1 Auxiliary Information
Depth Information serves as a crucial geometric cue for 3D scene reconstruction, particularly valuable in dynamic scenarios where it helps mitigate ambiguities in scale, motion, and geometry. Through gradient backpropagation, supervising the rendered depth in observation space significantly aids the learning of both radiance fields and motion fields [84, 197], particularly beneficial in challenging outdoor environments [96, 94, 98, 95]. When dedicated depth sensors are unavailable, monocular depth estimation methods [198, 199] can provide valuable geometric cues despite scale ambiguity [16, 168, 83, 197]. Beyond providing additional geometric supervision, depth information also enables importance sampling in regions near object surface, substantially reducing unnecessary computational overhead in volume-based rendering methods [200, 71].
@@
Surface Normals capture fine-grained geometric details that might be missed in neural representations, making them valuable for high-fidelity reconstruction. In dynamic scenes, surface normals are particularly useful in tracking shape deformations [201, 147] and modeling view-dependent reflections [202]. Recent research has successfully leveraged normal information to recover detailed human surface geometry [203, 204, 205], and enhance scene reconstruction [206, 207, 208]. For supervision purpose, pseudo ground truth normals can be derived using foundation models like Metric3D [209] or calculated from template meshes [36, 149, 158].
@@
Semantic Information serves as a powerful auxiliary cue for accurate modeling of moving objects. In dynamic scenes, object-level semantic segmentation provides valuable silhouettes or foreground masks that help localize dynamic objects and decompose scenes into static backgrounds and dynamic foregrounds [210, 202, 84]. This semantic decomposition is particularly valuable for complex urban environments with multiple moving entities [22, 94]. Beyond object-level segmentation, part-level semantic information recognizes the rigid components of articulated objects [211], enforcing part consistency throughout the dynamic reconstruction process and enabling more detailed part-wise reconstruction [212]. At the finest granularity, pixel-level semantic features facilitate robust point tracking and establish reliable correspondences across frames [213, 121, 172, 214].
@@
Data-driven Priors significantly enhance dynamic scene reconstruction by leveraging implicit knowledge from large-scale datasets to provide valuable constraints during optimization. For example, optical flow models like UniMatch [215] establish dense correspondences between frames [119, 151, 152, 187], offering important cues for flow field supervision, while object tracking models generate reliable trajectories for rigid objects in motion. Additionally, visual foundation models like DINO [153] extract semantic features that maintain consistency across both spatial viewpoints and temporal frames, providing robust cues for establishing correspondences in challenging scenarios [188, 156, 120, 179]. These data-driven approaches substantially enhance reconstruction fidelity in complex and dynamic environments where manual annotations remain prohibitively labor-intensive to acquire.
@@ 4.5.2 Regularization
Physical constraints play a crucial role in dynamic scene reconstruction by enforcing motion continuity and structural preservation. Temporal and spatial smoothness is achieved through Total Variation (TV) loss, which encourages piecewise constant motion with natural transitions between different regions [81, 14, 216], while Laplacian regularization penalizes sharp gradient changes in deformation fields [143]. Structural integrity during deformation is maintained through As-Rigid-As-Possible (ARAP) constraints, which preserve local neighborhood relationships, and isometric loss, which maintains global geodesic distances. Additionally, divergence loss promotes volume preservation by constraining deformations to primarily consist of translations and rotations, a critical property for realistic object and scene modeling.
Scene priors further enhance reconstruction quality by incorporating domain knowledge into the optimization process. $L_{1}$ regularization on motion fields biases scenes toward static components [26, 183], while cycle consistency enforces coherent correspondences between features, geometry, and appearance across frames [38, 13, 188]. For visual quality improvement, opacity regularization promotes binary outcomes (0 or 1) to eliminate floating artifacts [26, 183]. These complementary regularizations collectively ensure that dynamic scene reconstructions achieve physical accuracy, temporal consistency, and visual plausibility.
Figure 5: We propose a unified framework to encapsulate various representation paradigms.